Let X = {a, b, c} defined by f : X ®X such that f = {(a, b), (b, a), (c,c)}.  Find the values of

f–1, f2 and f4


1
Expert's answer
2021-09-20T16:11:55-0400

Let "X = \{a, b, c\}" defined by "f : X \to X" such that "f = \{(a, b), (b, a), (c,c)\}" .

Let us find the values of "f^{\u20131}, f^2" and "f^4". Since "f^{-1}(y)=x" iff "f(x)=y," we concluse that "f^{-1} = \{(b, a), (a, b), (c,c)\}." Taking itno aaccount that "f^2(x)=f(f(x))" and "f^4(x)=f^2(f^2(x))," we conclude that "f^2 = \{(a, a), (b, b), (c,c)\}" and "f^4 = \{(a, a), (b, b), (c,c)\}."